# Investigating Adversarial Attacks in Software Analytics via Machine Learning Explainability

### v1

@@ 3.1.2. Feature Engineering
The rationale for selecting six datasets in this paper is described in Section 3.1.1. Therefore, to optimize the performance of ML models in our experiment, we adhere to the same feature engineering techniques previously applied to these datasets by other studies. In the study conducted by Catolino et al. (Catolino et al., 2019), the InfoGain method was employed to select a subset of features. From the original 14 features, this technique identified 6 as particularly relevant, which we incorporated into our model. Similarly, Nafi et al. (Nafi et al., 2019) conducted feature selection for their code fragments, retaining 9 features from an initial pool of 24. This process resulted in the derivation of 18 feature values, each associated with a single class label for every clone pair. We utilized these filtered features to train ML models in our study. When dealing with the Java project dataset, we encounter the issue of collinearity among features. To address this challenge, we apply the AutoSpearman technique, which has been utilized by Yatish et al. (Yatish et al., 2019) on the same dataset. Applying this technique allows us to identify and retain 27 pertinent features from an initial set of 65. In our investigations on the Postgres dataset, we also encounter the challenge of the multi-collinearity problem. To address this issue effectively, we implement the same feature selection and normalization methodologies outlined in Kamei et al. (Kamei et al., 2012) on this dataset.
@@
Feng et al. (Feng et al., 2024) proposed a code clone detection method, namely Toma, using the BCB dataset and classical machine learning models. They converted the source code into type sequences and applied six similarity metrics, such as the Jaccard similarity coefficient, Dice similarity coefficient, Levenshtein distance, Levenshtein ratio, Jaro similarity, and Jaro-Winkler similarity, to extract six features from the type sequences. Thus, the BCB dataset has six features and is used in our study to train the ML models.
3.1.3. Dataset Balancing
@@
The JIT defect prediction task faces the challenge of imbalanced binary classification, with one class (e.g., clean commit) having significantly more examples than the other (e.g., buggy commit). To tackle this issue, we employ the Synthetic Minority Over-sampling Technique (SMOTE), a technique previously utilized in previous studies (Kamei et al., 2012; Roy et al., 2022; Pornprasit et al., 2021) on the defect prediction datasets selected for our study. Additionally, we apply SMOTE to rectify imbalanced datasets within the cross-language clone dataset. We must emphasize that we exclusively use this technique for the training data, ensuring that the test dataset remains unaffected by the oversampling process. Finally, the BCB dataset has more clone pairs than non-clone pairs. Since the BCB dataset contains only 278,838 non-clone pairs, we randomly select approximately 270,000 clone pairs from eight million clones, following the study by Feng et al. (Feng et al., 2024), ensuring that the BCB dataset used in our study is balanced.
3.2. Machine Learning Model Selection and Training
@@
Selecting and training ML models are crucial for the success of any ML project. This process entails choosing the appropriate algorithm or model architecture, training it on the preprocessed dataset, and tuning hyperparameter values to enhance generalization while avoiding overfitting.
3.2.1. Machine Learning Models
@@
Ravid et al. (Shwartz-Ziv and Armon, 2022) demonstrated that classical ML models trained on tabular data outperformed DL models trained on the same data in various tasks. In addition, Feng et al. (Feng et al., 2024), Hu et al. (Hu et al., 2022), and Wu et al. (Wu et al., 2022) demonstrated that ML models (e.g., Random Forest) outperformed several state-of-the-art deep learning models ((e.g., ASTNN (Zhang et al., 2019), SCDetector (Wu et al., 2020), DeeSIM (Zhao and Huang, 2018), FCCA (Hua et al., 2020), CDLH (Wei and Li, 2017), and TBCNN (Mou et al., 2014))) in software clone detection on the BCB dataset. Since the datasets selected for this study contain tabular data, we aim to evaluate the robustness of ML models trained on features extracted from the original input (e.g., source code) presented in a tabular format. Consequently, we focus on ML models commonly used for tabular data in various software analytics tasks, including defect prediction, clone detection, and classification of code review comments. For example, Logistic Regression and Random Forest are commonly used classification techniques in software defect prediction. Thus, we select widely used ML models for our experiments, encompassing three standard ML models (Logistic Regression (LR), Multi-Layered Perceptron (MLP), and Decision Tree (DT)), as well as four ensemble methods (Gradient Boosting Classifier (GBC), Random Forest (RF), AdaBoost (ADA), and Bagging (BAG)). It is worth noting that all of these models have been successfully employed in previous software analytics studies (Awal and Roy, 2024; Kamei et al., 2012; Pornprasit et al., 2021; Catolino et al., 2019; Roy et al., 2022; Rahman et al., 2017; Feng et al., 2024; Hu et al., 2022; Wu et al., 2022).
3.2.2. Hyperparameters Tuning
@@
In the realm of model optimization, the tuning of hyperparameters plays a pivotal role in facilitating effective generalization and mitigating the risk of overfitting. To maximize model performance, we explore various hyperparameter combinations, following established standard settings for different ML models (Kamei et al., 2012; Pornprasit et al., 2021; Catolino et al., 2019; Roy et al., 2022; Rahman et al., 2017; Tantithamthavorn et al., 2016, 2018; Feng et al., 2024). For example, Feng et al. (Feng et al., 2024) observed that setting the depth parameters to 32, 32, 64, and 16 for RF, DT, Adaboost, and GBDT models, respectively, achieves the highest F1-scores for clone detection on the BCB dataset. In addition, our approach leverages various optimization techniques, including grid search (Liashchynskyi and Liashchynskyi, 2019), random search (Bergstra and Bengio, 2012), and Bayesian optimization (Wu et al., 2019), to meticulously fine-tune hyperparameters, ultimately selecting the optimal combinations for our specific objectives. This rigorous optimization process is a cornerstone of our ML model training, ensuring that our models are finely calibrated to yield the best possible performance. It is important to note that we only tune hyperparameters while leaving the model’s parameters unaltered.
3.2.3. Evaluate Model Performance
@@
Model validation is a critical step in ML, ensuring the model’s ability to perform well on new, unseen data. Since PyExplainer often requires excessive time to generate an explanation for a single instance, we adopt a dataset-splitting approach, dividing it into training (90%) and testing (10%) subsets to expedite the overall process. To validate the ML models using the 90% training data, we employ a 10-fold cross-validation technique, which was applied in other software analytics studies (Awal and Roy, 2024; Feng et al., 2024; Hu et al., 2022; Wu et al., 2022; Roy et al., 2022; Kamei et al., 2012; Yatish et al., 2019). Finally, we conduct a thorough evaluation of model performance, utilizing key metrics, including accuracy, F1-score, and the Area Under the Receiver Operating Characteristic Curve (AUC), as recommended by recent studies (Yatish et al., 2019; Catolino et al., 2019; Roy et al., 2022).
@@
The AUC assesses the discriminatory power of models by measuring the true negative rate (coverage of the negative class) on the $x$-axis and the true positive rate (coverage of the positive class) on the $y$-axis. AUC values range from 00 (indicating the worst performance) to $0.5$ (no better than random guessing) up to $1$ (representing the best performance) (Hanley and McNeil, 1982). Therefore, AUC-ROC is a comprehensive tool for evaluating the effectiveness of binary classification models in distinguishing between positive instances (e.g., buggy commit) and negative instances (e.g., clean commit). Finally, since we aim to perform explanation-guided adversarial attacks on ML models, a test-AUC value of 0.75 is the minimum threshold for ensuring the reliability of explanations (Roy et al., 2022; Lyu et al., 2021). Thus, we evaluate the performance of ML models trained on the selected datasets for this study.
3.3. ML Explainability Technique Selection
@@
This study aims to investigate adversarial attacks in software analytics tasks using the ML explainability technique. Our goal is to modify the fewest features possible when generating adversarial examples (Mathov et al., 2020). Therefore, ML explainability techniques can serve as an effective solution for identifying the important features on which ML models base their decisions. We then modify the top-$k$ features from the important feature list to generate adversarial examples, where $k$ ranges between 1 and half of the total features. The details of adversarial attacks using ML explainability techniques are described in Section 3.4.
@@
While numerous ML explainability techniques have been proposed in recent years to shed light on the decisions made by ML models (Lundberg and Lee, 2017; Ribeiro et al., 2016; Tantithamthavorn and Jiarpakdee, 2021), our focus is solely on model-agnostic techniques, which are applied across various software analytics tasks. Jiarpakdee et al. (Jiarpakdee et al., 2020) empirically demonstrated the effectiveness of model-agnostic ML explainability techniques (e.g., LIME (Ribeiro et al., 2016)) in elucidating the outcomes of defect prediction models. Recently, Roy et al. (Roy et al., 2022) demonstrated that model-agnostic techniques such as SHAP and LIME could be successfully used for post-hoc analysis of black-box ML models in software analytics tasks. Additionally, Amich demonstrated that LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017) yield more accurate results when compared to other model-agnostic explanation methods, such as DeepLIFT (Shrikumar et al., 2017) and LEMNA (Guo et al., 2018). Therefore, we select two highly cited model-agnostic ML explainability techniques, SHAP and LIME. Furthermore, Pornprasit et al. (Pornprasit et al., 2021) introduced the model-agnostic technique PyExplainer, which builds upon the concept of LIME for JIT defect prediction models and supports all ML models available in the scikit-learn library. Hence, we also include PyExplainer in our study due to its model-agnostic nature.
3.4. Adversarial Attacks
@@
In computer vision (Szegedy et al., 2013), a carefully crafted pixel-level perturbation added to an input image yields an adversarial example that cannot be differentiated from the original input upon visual inspection. The main property of adversarial examples is that they are inputs intentionally modified to cause incorrect predictions while remaining imperceptible to the human eye. However, this is not true for any software analytics task. In software analytics, any transformations to feature values in tabular format or tokens of source code lead to perceptible changes to the human eye. Therefore, adversarial example generation in software analytics tasks differs fundamentally from that in computer vision. Ballet et al. (Ballet et al., 2019), and Cartella et al. (Cartella et al., 2021) introduced a distinct concept of imperceptibility concerning adversarial examples within financial domain data. They altered non-important features based on human judgment while generating these examples. For instance, according to expert analysis, only a subset of features, such as income and age, might be crucial for specific predictions (e.g., fraudulent transaction detection). The attacker should not aim to change these important features; instead, they should target the non-important features so that the generated adversarial examples remain imperceptible. However, these imperceptibility concepts are domain-specific and are not directly applicable to tabular data in software analytics tasks. Therefore, to ensure imperceptibility while attacking, we consider the criterion introduced by Mathov et al. (Mathov et al., 2020), which states that “an attacker should aim to minimize the number of modified features (e.g., a minimal $\ell_{0}$ perturbation) when dealing with tabular data.” The $\ell_{0}$ perturbation is defined as follows:
@@
Table 1. An instance that we use to generate an explanation in Fig. 2(c) and calculation of one standard deviation (STD) value of each feature using the whole training dataset.
@@
Figure 2. Important features found by SHAP (Fig. 2(a)), LIME (Fig. 2(b)), and PyExplainer (Fig. 2(c)) on the cross-project mobile apps dataset. Figure 2(c) depicts the explanation and Risk Score when PyExplainer works on the original input. Figure 2(d) shows the explanation and Risk Score when PyExplainer works on the modified features in the guided direction.
@@
To determine the value of $k$ for the top-k important features, we leveraged the concept of the Elbow method in the $K$-means clustering algorithm (Kodinariya et al., 2013), which we refer to as the Reverse Elbow Method. Initially, we modify the top-$1$ feature, proceed to the top-$2$ features, and so forth to calculate the $ASR$ (definition of ASR is provided in Section 3.5.) metric value until we reach half the total number of features in the feature importance rank. For example, in Figure 2(a), the feature importance rank of the cross-project mobile apps dataset is displayed from top to bottom. As a result, when considering half of the features (e.g., $3$ out of $6$ features), we obtain three feature combinations (Line $7$, Algorithm 1): {ndev}, {ndev, nd}, and {ndev, nd, nf}, as shown in Table 2.
Table 2. Determining optimal top-$k$ features for adversarial attacks using ASR metric values.
@@ 26                   $Top_{k}\leftarrow len(FC_{ASR}[i])$
Algorithm 1 Explanation-guided Adversarial Attacks
We leverage the technique introduced in the previous study by Pornprasit et al. (Pornprasit et al., 2021) to address the second aspect and perform the adversarial attacks (Lines 8–12, Algorithm 1), and this is one of the many ways to change the feature values (Hashemi and Fathi, 2020). PyExplainer explicitly identifies the features that influence ML models’ decision-making processes. It also offers threshold values for each feature, enabling us to make additions or subtractions of one STD with the threshold values in alignment with the guided directions for generating adversarial examples. We calculate STD from the training set as mentioned in the previous study (Pornprasit et al., 2021). As SHAP and LIME do not provide specific threshold values with their explanations, we perform one STD addition and subtraction operation relative to the original feature values. It is crucial to note that we discard changes that would result in a negative value after subtracting one STD from the current value. This precaution ensures that the transformed instances remain valid, do not violate the fundamental properties of the features, and do not generate out-of-distribution samples. For example, the feature number of modified directories (nd) can’t have a negative value after its alteration (e.g., $1-14=-13$).
@@
3.5. Robustness Evaluation
We evaluate the robustness of the selected ML models using the generated adversarial examples. To assess the quality of these adversarial examples and evaluate the robustness of the ML models, we employ the Attack Success Rate (ASR) metric, as previously used by Yang et al. (Yang et al., 2022). The ASR quantifies the percentage of instances that were correctly predicted but are no longer predicted correctly by the ML models following adversarial attacks. The ASR is defined as follows:
(2) $ASR=\frac{|\{x|x\in X\wedge M(x^{\prime})\neq M(x)\}|}{|X|}$

### v2

@@ 3.1.2 Feature Engineering
The rationale for selecting six datasets in this paper is described in Section 3.1.1. Therefore, to optimize the performance of ML models in our experiment, we adhere to the same feature engineering techniques previously applied to these datasets by other studies. In the study conducted by Catolino et al. [13], the InfoGain method was employed to select a subset of features. From the original 14 features, this technique identified 6 as particularly relevant, which we incorporated into our model. Similarly, Nafi et al. [1] conducted feature selection for their code fragments, retaining 9 features from an initial pool of 24. This process resulted in the derivation of 18 feature values, each associated with a single class label for every clone pair. We utilized these filtered features to train ML models in our study. When dealing with the Java project dataset, we encounter the issue of collinearity among features. To address this challenge, we apply the AutoSpearman technique, which has been utilized by Yatish et al. [14] on the same dataset. Applying this technique allows us to identify and retain 27 pertinent features from an initial set of 65. In our investigations on the Postgres dataset, we also encounter the challenge of the multi-collinearity problem. To address this issue effectively, we implement the same feature selection and normalization methodologies outlined in Kamei et al. [15] on this dataset.
@@
Feng et al. [2] proposed a code clone detection method, namely Toma, using the BCB dataset and classical machine learning models. They converted the source code into type sequences and applied six similarity metrics, such as the Jaccard similarity coefficient, Dice similarity coefficient, Levenshtein distance, Levenshtein ratio, Jaro similarity, and Jaro-Winkler similarity, to extract six features from the type sequences. Thus, the BCB dataset has six features and is used in our study to train the ML models.
3.1.3 Dataset Balancing
@@
The JIT defect prediction task faces the challenge of imbalanced binary classification, with one class (e.g., clean commit) having significantly more examples than the other (e.g., buggy commit). To tackle this issue, we employ the Synthetic Minority Over-sampling Technique (SMOTE), a technique previously utilized in previous studies [15, 65, 35, 66] on the defect prediction datasets selected for our study. Additionally, we apply SMOTE to rectify imbalanced datasets within the cross-language clone dataset. We must emphasize that we exclusively use this technique for the training data, ensuring that the test dataset remains unaffected by the oversampling process. Finally, the BCB dataset has more clone pairs than non-clone pairs. Since the BCB dataset contains only 278,838 non-clone pairs, we randomly select approximately 270,000 clone pairs from eight million clones, following the study by Feng et al. [2], ensuring that the BCB dataset used in our study is balanced.
3.2 Machine Learning Model Selection and Training
@@
Selecting and training ML models is crucial for the success of any ML project. This process entails choosing the appropriate algorithm or model architecture, training it on the preprocessed dataset, and tuning hyperparameter values to enhance generalization while avoiding overfitting.
3.2.1 Machine Learning Models
@@
Ravid et al. [51] demonstrated that classical ML models trained on tabular data outperformed DL models trained on the same data in various tasks. In addition, Feng et al. [2], Hu et al. [3], and Wu et al. [4] demonstrated that ML models (e.g., Random Forest) outperformed several state-of-the-art deep learning models (e.g., ASTNN [44], SCDetector [45], DeeSIM [46], FCCA [47], CDLH [48], and TBCNN [49]) in software clone detection on the BCB dataset. Since the datasets selected for this study contain tabular data, we aim to evaluate the robustness of classical ML models trained on features extracted from the original input (e.g., source code) presented in a tabular format. Consequently, we focus on ML models commonly used for tabular data in various software analytics tasks, including defect prediction, clone detection, and classification of code review comments. For example, Logistic Regression and Random Forest are commonly used classification techniques in software defect prediction [67, 68]. Thus, we select widely used ML models for our experiments, encompassing three standard ML models (Logistic Regression (LR), Multi-Layered Perceptron (MLP), and Decision Tree (DT)), as well as four ensemble methods (Gradient Boosting Classifier (GBC), Random Forest (RF), AdaBoost (ADA), and Bagging (BAG)). It is worth noting that all of these models have been successfully employed in previous software analytics studies [66, 15, 35, 13, 65, 39, 2, 3, 4, 67, 68].
3.2.2 Hyperparameters Tuning
@@
In the realm of model optimization, the tuning of hyperparameters plays a pivotal role in facilitating effective generalization and mitigating the risk of overfitting. To maximize model performance, we explore various hyperparameter combinations, following established standard settings for different ML models [15, 35, 13, 65, 39, 69, 70, 2]. For example, Feng et al. [2] observed that setting the depth parameters to 32, 32, 64, and 16 for RF, DT, Adaboost, and GBDT models, respectively, achieves the highest F1-scores for clone detection on the BCB dataset. In addition, our approach leverages various optimization techniques, including grid search [71], random search [72], and Bayesian optimization [73], to meticulously fine-tune hyperparameters, ultimately selecting the optimal combinations for our specific objectives. This rigorous optimization process is a cornerstone of our ML model training, ensuring that our models are finely calibrated to yield the best possible performance. It is important to note that we only tune hyperparameters while leaving the model’s parameters unaltered.
3.2.3 Evaluate Model Performance
@@
Model validation is a critical step in ML, ensuring the model’s ability to perform well on new, unseen data. Since PyExplainer often requires excessive time to generate an explanation for a single instance, we adopt a dataset-splitting approach, dividing it into training (90%) and testing (10%) subsets to expedite the overall process. To validate the ML models using the 90% training data, we employ a 10-fold cross-validation technique, which was applied in other software analytics studies [66, 2, 3, 4, 65, 15, 14]. Finally, we conduct a thorough evaluation of model performance, utilizing key metrics, including accuracy, F1-score, and the Area Under the Receiver Operating Characteristic Curve (AUC), as recommended by recent studies [14, 13, 65].
@@
The AUC assesses the discriminatory power of models by measuring the true negative rate (coverage of the negative class) on the $x$-axis and the true positive rate (coverage of the positive class) on the $y$-axis. AUC values range from 00 (indicating the worst performance) to $0.5$ (no better than random guessing) up to $1$ (representing the best performance) [74]. Therefore, AUC-ROC is a comprehensive tool for evaluating the effectiveness of binary classification models in distinguishing between positive instances (e.g., buggy commit) and negative instances (e.g., clean commit). Finally, since we aim to perform explanation-guided adversarial attacks on ML models, a test-AUC value of 0.75 is the minimum threshold for ensuring the reliability of explanations [65, 75]. Thus, we evaluate the performance of ML models trained on the selected datasets for this study.
3.3 ML Explainability Technique Selection
@@
This study aims to investigate adversarial attacks in software analytics tasks using the ML explainability technique. Our goal is to modify the fewest features possible when generating adversarial examples [76]. Therefore, ML explainability techniques can serve as an effective solution for identifying the important features on which ML models base their decisions. We then modify the top-$k$ features from the important feature list to generate adversarial examples, where $k$ ranges between 1 and half of the total features. The details of adversarial attacks using ML explainability techniques are described in Section 3.4.
@@
While numerous ML explainability techniques have been proposed in recent years to shed light on the decisions made by ML models [57, 58, 77], our focus is solely on model-agnostic techniques, which are applied across various software analytics tasks. Jiarpakdee et al. [36] empirically demonstrated the effectiveness of model-agnostic ML explainability techniques (e.g., LIME [58]) in elucidating the outcomes of defect prediction models. Recently, Roy et al. [65] demonstrated that model-agnostic techniques such as SHAP and LIME could be successfully used for post-hoc analysis of black-box ML models in software analytics tasks. Additionally, Amich demonstrated that LIME [58] and SHAP [57] yield more accurate results when compared to other model-agnostic explanation methods, such as DeepLIFT [78] and LEMNA [79]. Therefore, we select two highly cited model-agnostic ML explainability techniques, SHAP and LIME. Furthermore, Pornprasit et al. [35] introduced the model-agnostic technique PyExplainer, which builds upon the concept of LIME for JIT defect prediction models and supports all ML models available in the scikit-learn library. Hence, we also include PyExplainer in our study due to its model-agnostic nature.
3.4 Adversarial Attacks
@@
In computer vision [18], a carefully crafted pixel-level perturbation added to an input image yields an adversarial example that cannot be differentiated from the original input upon visual inspection. The main property of adversarial examples is that they are inputs intentionally modified to cause incorrect predictions while remaining imperceptible to the human eye. However, this is not true for any software analytics task. In software analytics, any transformations to feature values in tabular format or tokens of source code lead to perceptible changes to the human eye. Therefore, adversarial example generation in software analytics tasks differs fundamentally from that in computer vision. Ballet et al. [80], and Cartella et al. [81] introduced a distinct concept of imperceptibility concerning adversarial examples within financial domain data. They altered non-important features based on human judgment while generating these examples. For instance, according to expert analysis, only a subset of features, such as income and age, might be crucial for specific predictions (e.g., fraudulent transaction detection). The attacker should not aim to change these important features; instead, they should target the non-important features so that the generated adversarial examples remain imperceptible. However, these imperceptibility concepts are domain-specific and are not directly applicable to tabular data in software analytics tasks. Therefore, to ensure imperceptibility while attacking, we consider the criterion introduced by Mathov et al. [76], which states that “an attacker should aim to minimize the number of modified features (e.g., a minimal $\ell_{0}$ perturbation) when dealing with tabular data.” The $\ell_{0}$ perturbation is defined as follows:
@@
Table 2: An instance that we use to generate an explanation in Fig. 2(c) and calculation of one standard deviation (STD) value of each feature using the whole training dataset
@@
Figure 2: Important features found by SHAP (Fig. 2(a)), LIME (Fig. 2(b)), and PyExplainer (Fig. 2(c)) on the cross-project mobile apps dataset. Figure 2(c) depicts the explanation and Risk Score when PyExplainer works on the original input. Figure 2(d) shows the explanation and Risk Score when PyExplainer works on the modified features in the guided direction.
@@
To determine the value of $k$ for the top-k important features, we leveraged the concept of the Elbow method in the $K$-means clustering algorithm [82], which we refer to as the Reverse Elbow Method. Initially, we modify the top-$1$ feature, proceed to the top-$2$ features, and so forth to calculate the %ASR (definition of %ASR [^3] In this study, %ASR and ASR represent the same metric. is provided in Section 3.5.) metric value until we reach half the total number of features in the feature importance rank. For example, in Figure 2(a), the feature importance rank of the cross-project mobile apps dataset is displayed from top to bottom. As a result, when considering half of the features (e.g., $3$ out of $6$ features), we obtain three feature combinations (Line $7$, Algorithm 1): {ndev}, {ndev, nd}, and {ndev, nd, nf}, as shown in Table 3.
Table 3: Determining optimal top-$k$ features for adversarial attacks using ASR metric values
@@ 26                   $Top_{k}\leftarrow len(FC_{ASR}[i])$
Algorithm 1 Explanation-guided Adversarial Attacks.
We leverage the technique introduced in the previous study by Pornprasit et al. [35] to address the second aspect and perform the adversarial attacks (Lines 8–12, Algorithm 1), and this is one of the many ways to change the feature values [61]. PyExplainer explicitly identifies the features that influence ML models’ decision-making processes. It also offers threshold values for each feature, enabling us to make additions or subtractions of one STD with the threshold values in alignment with the guided directions for generating adversarial examples. We calculate STD from the training set as mentioned in the previous study [35]. As SHAP and LIME do not provide specific threshold values with their explanations, we perform one STD addition and subtraction operation relative to the original feature values. It is crucial to note that we discard changes that would result in a negative value after subtracting one STD from the current value. This precaution ensures that the transformed instances remain valid, do not violate the fundamental properties of the features, and do not generate out-of-distribution samples. For example, the feature number of modified directories (nd) can’t have a negative value after its alteration (e.g., $1-14=-13$).
@@
3.5 Robustness Evaluation
We evaluate the robustness of the selected ML models using the generated adversarial examples. To assess the quality of these adversarial examples and evaluate the robustness of the ML models, we employ the Attack Success Rate (ASR) metric, as previously used by Yang et al. [25]. The ASR quantifies the percentage of instances that were correctly predicted but are no longer predicted correctly by the ML models following adversarial attacks. The ASR is defined as follows:
%$\%ASR=\frac{|\{x|x\in X\wedge M(x^{\prime})\neq M(x)\}|}{|X|}$ (2)